OpenStudy (kittiwitti1):

When solving a simple trigonometric equation, first isolate the trigonometric function, and then determine the values for the _________ itself...? *adding another in progress*

OpenStudy (freckles):

I guess they are looking for variable... that is how I would solve something like tan(theta)+1=0 isolate the trig function tan(theta)=-1 then find theta such that the equation is true...

OpenStudy (kittiwitti1):

Okay - internet is being almighty lag-monster today, sorry -

OpenStudy (kittiwitti1):

That's not an option, these are the options http://prntscr.com/987w9m

OpenStudy (freckles):

only one of those make sense

OpenStudy (freckles):

what is your guess from the choices

OpenStudy (freckles):

If I asked you to evaluate tan(45 deg) what is 45 deg what does degrees measure ?

OpenStudy (kittiwitti1):

"angles" is wrong

OpenStudy (freckles):

oh that was the only option that made sense :p

OpenStudy (kittiwitti1):

submitted that before and was told it was incorrect, so I put this up here since I wouldn't know any other answer besides y-coordinate and trig function (both of which were marked wrong apparently)

OpenStudy (freckles):

ok well you said there was another question?

OpenStudy (freckles):

and this was the exact thing that was on your thingy right: "When solving a simple trigonometric equation, first isolate the trigonometric function, and then determine the values for the _________ itself...? "

OpenStudy (kittiwitti1):

sorry OS literally lag-crashed

OpenStudy (kittiwitti1):

There is but my internet's being laggy and OS is worse haha

OpenStudy (kittiwitti1):

http://prntscr.com/987xxt

OpenStudy (freckles):

well isolate the trig function first... \[\sin(\theta)=\frac{\sqrt{2}}{2}\] then whip out your unit circle and find when sin(theta) is sqrt(2)/2 also sin(theta) is the y-coordinate on the unit circle

OpenStudy (freckles):

so you are scanning the unit circle for the pairs (x,y) where y is sqrt(2)/2

OpenStudy (freckles):

and giving me the angle that corresponds to that y

OpenStudy (kittiwitti1):

ALMIGHTY LAG what gimme a sec to read that lol

OpenStudy (kittiwitti1):

Oh, it can work like that?

OpenStudy (kittiwitti1):

0-0

OpenStudy (freckles):

are you having trouble?

OpenStudy (freckles):

there is two values in (0,360) one is in the first quadrant the other is in the 2nd quadrant such that y is sqrt(2)/2

OpenStudy (kittiwitti1):

45 degrees and umm 180+45

OpenStudy (kittiwitti1):

180-45*

OpenStudy (freckles):

45 deg is right

OpenStudy (kittiwitti1):

135?

OpenStudy (freckles):

and 180-45 or 135 is right

OpenStudy (kittiwitti1):

Yay :]

OpenStudy (freckles):

so sin(theta)=sqrt(2)/2 when theta=45 or 135 deg

OpenStudy (kittiwitti1):

this is part b right

OpenStudy (freckles):

then you just add +360*k on to both of those and you have all your solutions

OpenStudy (kittiwitti1):

oh okay

OpenStudy (freckles):

since sin has period 360 deg

OpenStudy (kittiwitti1):

http://prntscr.com/987zyo

OpenStudy (kittiwitti1):

???

OpenStudy (kittiwitti1):

I'm going to try this but I don't know if it will work http://prntscr.com/98804t

OpenStudy (freckles):

you know it says missing deg symbol right? and also you didn't put the other solution to the first one

OpenStudy (kittiwitti1):

?

OpenStudy (kittiwitti1):

oh 135?

OpenStudy (freckles):

remember we got 45 and 135 for the second one and all it is saying is that it wants you to put the deg symbol for that and then for the first one I asked you to put +360k on to both so you should have 45+360k and 135+360k and yes since it is complaining about a deg symbol you will need to put one on 45 and 135 and 360

OpenStudy (kittiwitti1):

okay lol

OpenStudy (freckles):

or you can put it around (45+360k) and around (135+360k)

OpenStudy (kittiwitti1):

it just took off a point for the first question I posted what the heck.

OpenStudy (kittiwitti1):

either my internet lagged or the site is stupid

OpenStudy (kittiwitti1):

http://prntscr.com/98819l

OpenStudy (freckles):

isolate the trig function

OpenStudy (freckles):

and do exactly what we did for the other one

OpenStudy (kittiwitti1):

yes but sin(theta)-3 I don't know how to solve

OpenStudy (kittiwitti1):

OH WAIT \[\frac{\sqrt{3}}{2}\]

OpenStudy (kittiwitti1):

wow that was dumb of me

OpenStudy (freckles):

do you know how to solve \[2y-\sqrt{3}=0 \text{ for } y?\] yes add sqrt(3) on both sides then divide both sides by 2

OpenStudy (freckles):

\[\sin(\theta)=\frac{\sqrt{3}}{2}\]

OpenStudy (freckles):

now again y is the sine value

OpenStudy (kittiwitti1):

30+360k, 150+360k?

OpenStudy (freckles):

those are both incorrect

OpenStudy (freckles):

when theta is 30 y is 1/2 when theta is 150 y is 1/2

OpenStudy (freckles):

when theta is ____ y is sqrt(3)/2 and when theta is ____ y is sqrt(3)/2 y is the second number in the ordered pair

OpenStudy (kittiwitti1):

ugh I messed up pi/3 with ppi/6 sorry

OpenStudy (kittiwitti1):

pi*

OpenStudy (freckles):

so you mean 60 and 120?

OpenStudy (kittiwitti1):

yes

OpenStudy (freckles):

that would be correct

OpenStudy (freckles):

and please we learned from the last question it wants you to put in those degree symbols

OpenStudy (kittiwitti1):

lol ik

OpenStudy (freckles):

anyways peace and good luck

OpenStudy (freckles):

sleepy time

OpenStudy (kittiwitti1):

alright thank you and good night :]